OpenStudy (anonymous):

The wavelength (w) of a wave varies inversely as its frequency (f). A wave with a frequency of 600 kHz has a length of 207 m. What is the length of a wave with a frequency of 900 kHz? Length of the wave is m=

OpenStudy (anonymous):

w = k/f, 207m = k/600{kHz}, k= 1242 {kHz*m}, w = 1242{kHz*m} / 900{kHz} = 285.66 {m}

OpenStudy (anonymous):

so 285.66m?

OpenStudy (anonymous):

Yes